Find the missing values that solve this equation.
y = 5 -4x
(5, ____) (1, ____) (-1, ____)

Answer by funmath(2933)

Find the missing values that solve this equation.
y = 5 -4x
(5, ____) (1, ____) (-1, ____)
Ordered pairs are in this order: (x,y)
:
For (5,_) let x=5 and solve for y:
y=5-4(5)
y=5-20
y=-15
The ordered pair is: (5,-15)
:
For (1,_), let x=1 and solve for y:
y=5-4(1)
y=5-4
y=1
The ordered pair is (1,1)
:
For (-1,_), let x=-1 and solve for y:
y=5-4(-1)
y=5+4
y=9
The ordered pair is (-1,9)
:
For fun find (2,_) and (-2,_)
You should get (2,-3) and (-2,13) if you're doing it right.
Happy Calculating!!!